find two consecutive intergers such that twice the first is 26 more than the second
Answer by stanbon(75887)

find two consecutive integers such that
twice the first is 26 more than the second
----
1st: x
2nd: x+1
----
Equation:
2x = (x+1)+26
---
2x = x + 27
x = 27 (1st)
x+1 = 28 (2nd)
===================
Cheers,
Stan H.
===========